What is the probability that the first offspring from this cross:

Aa bb Dd Ee x AA bb Dd Ee

will be a son with the genotype:

Aa bb DD ee

Type your answer as a decimal and round to 6 decimal places.
Answer #1

The probability of Aa 1/2, DD 1/4, ee 1/4.

Hence 1/2*1/4*1/4 = 1/32

1/32 = 0.031250